# WavLLM: Towards Robust and Adaptive Speech Large Language Model

### v2

@@ 5 Conclusion and Future Work
In this paper, we propose WavLLM, a robust and adaptive speech large language model, which uses LLaMA-2-chat as foundational LLM backbone, and extracts semantic and acoustic information from speech audio utilizing Whisper and WavLM encoders. Utilizing a curriculum learning approach, the proposed WavLLM commences with single-task instruction training in the initial mixed training phase and subsequently expanding our training to encompass additional multi-task prompts in the second advanced phase with the integration of the proposed prompt adapter. Massive experiments demonstrate that our WavLLM model delivers state-of-the-art performance on various speech-related tasks and robust generalization abilities on single-task and multi-task instructions, enhanced by a CoT processing ability that greatly improves its effectiveness in tackling intricate tasks. For future work, we are interested in advancing the capability to autonomously decompose complex one-shot instruction tasks into CoT based tasks, and equipping our model with the speech synthesis ability.
Appendix A Some Examples of Training Data
A.1 SQA Task
speech transcript instructions targets source
the nautilus would still have nothing to fear from their attacks What animal would still have nothing to fear from attacks? The nautilus. LibriSpeech
as gelatinous or glutinous matters impress the sense of touch i have spoken both of sound and of voice i mean to say that the sound was one of distinct of even wonderfully thrillingly distinct What is the gender of the speaker? And what is the speaker talking about? The speaker is female. And she is talking about sound and voice being distinct. LibriSpeech
A: so how do you feel about it; B: uh i don’t see any problem with with random- testing for random drugs; A: yeah i myself almost um am in favor of it uh; A: i work at Honeywell and i went through a a pretty i don’t know i went through a standard drug testing thing before i was brought on; B: um-hum What is the second speaker’s opinion on random drug testing? The second speaker doesn’t see any problem with random drug testing. Switchboard
A.2 Multi-task Instruction Datasets
speech transcript instructions targets source
[…]
A.3 The Prompts for Generating SQA Data by GPT-4
You are asked to generate **only one** questions, and their corresponding answers, according to some conversational sentences given below. These sentences have been transcribed from conversational speech data with one or multiple speakers who are taking to each other. “Speaker A” and “Speaker B” in the senteces are labeled by human and your response must not contain human-marked information, namely “Speaker A” and “Speaker B”. Here are the requirements: 1. Your response should strictly follow the format below: “Question”: “xxx”, “Answer”: “xxx”; 2. Please ignore “Speaker A” and “Speaker B” in the given sentences. Your response should strictly not include the phrase “Speaker A” and “Speaker B”; 3. Your question should be highly related to the conversation, and your answer must be **correct**, and should be simple and clear. Besides, you question should be designed as your answer has to be reasoned from the conversation; 4. For example, a sentence “Speaker A: It is a good day; Speaker B: Yes, but I have to go to hospital” means that speaker A first say it is a good day and speaker B then say that Yes, but I have to go to hospital. 5. **Very Importance**: Your questions and answers **can not** contain the word “Speaker A” and “Speaker B”, because “Speaker A” and “Speaker B” in the sentences are additional labels for transcripts, and they are different people. For example, the question “What is Speaker B’s opinion?” **does not** meet the requirements because it contains word “Speaker B”. Namely, you can not use “Speaker A” and “Speaker B” to represent they in questions and answers, maybe you can use the first or second speaker to denote “Speaker A” or “Speaker B”; 6. The type of response should be diverse. The respone **must contain** double quotation marks for each part. Here are the sentences: transcript
Appendix B Some Examples of Evaluation Data
B.1 Examples of II-task Instruction
speech transcript instructions targets source
Women: “How much time do you usually spend exercising daily?” Man: “Frankly speaking, I’m an awfully lazy man. I know it’s time to change.” To begin, What will the man do next? A. Start to take exercise; B. Do as he always does; C. Change his working time.; Next, Create a French transcript for this English audio clip; Furthermore, Recognize the speech and give me the transcription; Last step, setting aside the audio, Who wrote “The Adventures of Sherlock Holmes”? - MuTual
B.2 Examples of CoT-task and Non-CoT-task Instruction
speech transcript instructions targets source
three films from Asia-Pacific are in the running for the coveted golden palms at this yearś Cannes film festival, competing in a field dominated by European productions, organizers announced Monday. First of all, transcribe the audio recording into text, capturing every spoken word; Additionally given this audio clip and text, can you condense it into a clear, concise summary, no more than 20 words?; Lastly disregarding the sound, translate this English summary into German. Drei Filme aus dem asiatisch-pazifischen Raum im Rennen in Cannes gigaword
three films from Asia-Pacific are in the running for the coveted golden palms at this yearś Cannes film festival, competing in a field dominated by European productions, organizers announced Monday. Please summarize the content of the audio clip in German, no more than 20 words. Drei Filme aus dem asiatisch-pazifischen Raum im Rennen in Cannes gigaword
Appendix C The Prompt for Scoring using GPT-4
C.1 SQA Scoring
Next, I will give you a multiple-choice question along with its correct answer, as well as a generated answer that needs to be evaluated for correctness. You will need to determine whether the given answer is correct based on the question and the correct answer, and give a simple reason. The answer must explicitly give the correct option to be considered correct and not by implication or indirect response. Your response should strictly follow the format:{"result": "xx", "reason": "xx"}, if the given answer is correct, then your response should be {"result": "True", "reason": "xx"}, otherwise your response should be {"result": "False", "reason": "xx"}.Here is the question: {"What will the man do next? A. Start to take exercise; B. Do as he always does; C. Change his working time."},and the correct answer is {"A"},the answer that needs to be judged is {"B. Do as he always does"}.
C.2 II-task Scoring
Next i will give you an audio transcription, instructions related or unrelated to the audio, and the corresponding responses. You need to use the given information to figure out how many instructions were completed correctly in the given responses and how many were left unanswered or answered incorrectly, then give the simple and clear reason why each question was completed or not, Finally, you need to do the self-examination to ensure responses you give must be correct and without inconsistencies.
You must adhere to the following rules: 1. instructions are only answered sequentially, and one answer must only correspond to one instruction; 2. For choice questions, the answer must be in **English** and only **one** of the three options, multiple options or none of the options are considered incomplete, e.g. the answer is “A. xx; B. xx; C. xx.” (note that options may appear across sentences or lines); 3. All instructions must be completed in **English**, except translation tasks; 4. The translation task is not a translation of instructions, but a translation of the content of the given audio speech 5. If the response is a repetition of the instruction, then it is considered not completed.
Your response should strictly follow the format: {“instructions completed”: “xx”, “instructions incompleted”: “xx”, “Reasons for each instructions”: [“instruction_1: xx”, “instruction_2: xx”, “instruction_3: xx”, “instruction_4: xx”]}.
You need to refer to this example: the transcription of the audio: “So James What are you going to do for your birthday? Well, I was hoping to have a party but most of my friends are busy so I have accepted my sisters offer to take me to a film. At least it is better than going out for a meal with my parents.”; the instructions are “To begin, ignore the audio clip, Who is known for developing the theory of relativity?; Then, What will the man do on his birthday? A. Have a party; B. See a movie; C. Go out for a meal.; Moreover, Convert the spoken words in this audio file into a textual format; Last step, Translate this English speech into Japanese";
and the respones are “1. {CJK*}UTF8min エイトリングの理論を発展した人物は誰ですか? ; 2. What will the man do on his birthday? A. Have a party; B. See a movie; C. Go out for a meal.; 3. {CJK*}UTF8minこの音声>ファイルをテキスト形式に変換してください。; 4. {CJK*}UTF8minこの英語のスピーチを日本語に翻訳してください。”.
And your output should be {“instructions completed”: “0”, “instructions incompleted”: “4”, “Reasons for each instructions”: [“instruction_1: incompleted, the answer is the Japanese translation of the instruction. Q&A task should be completed in English.”, “instruction_2: incompleted, the answer is the a repetition of the instruction. More than one options were given”, “instruction_3: incompleted, the answer is the Japanese translation of the instruction”, “instruction_4: incompleted, the answer is the Japanese translation of the instruction”]}.
Be sure to carefully follow the rules and refer to the examples for responses.
Here is the transcription of the audio: {“I really want to go to the concert, but I haven’t got a ticket yet leave it to me, I’ll organize that we can go together really that’s wonderful.”},
the instructions are {“To begin, disregarding the sound, How many elements are on the periodic table?; Then, Translate this English speech into Japanese; Additionally, How will the speaker get a ticket to the concert? A. The man will go to buy the ticket; B. The woman will get the ticket; C. The man will have someone buy the ticket.; Lastly, Transcribe the given audio clip into text”},
the responses is {“Addressing the prompt ’How many elements are on the periodic table?’, the answer is: There are currently 118 elements on the periodic table. From the audio, we have the following translations: {CJK*}UTF8minコンサートに行きたいけど まだチケットがないの どうするんだ? For the question “How will the speaker get a ticket to the concert? A. The man will go to buy the ticket; B. The woman will get the ticket; C. The man will have someone buy the ticket.”, the response is: B; The audio transcription is: i really want to go to the concert but i haven’t got a ticket yet b leave it to me i’ll organize that we can go together r really that’s wonderful }
Appendix D Prompt Template for WavLLM
[…]
Appendix E Example of Multi-round Dialog
Figure 3: An example of multi-round dialog based on Gaokao task
Appendix F Training Data Details
Training data used in the first stage and second stage. For all tasks, the instructions are diverse. “#Hours” refers to the duration of speech data for each task, not the total number of hours of the data source. The targets of SQA tasks are generated using GPT3.5, GPT-4 or LLaMA-2-chat. Task Description Data Source #Hours Single -task automatic speech recognition (ASR) LibriSpeech 960 LibriHeavy medium 1800 speech-to-text translation (ST), including English to German (En2De), English to Japanese (En2Ja), and English to Chinese (En2Zh) CoVoST2 440 MuST-C 280 speaker verification (SV) VoxCeleb 1290 emotion recognition (ER) IEMOCAP Session 1-4 5 speech question answering (SQA), including gender and speaker-related questions, and multi-round QA LibriSpeech 520 AMI 50 Fisher 710 Switchboard 230 speech question answering (SQA), continue writing tasks LibriSpeech 960 speech question answering (SQA), summary tasks LibriSpeech 410 instruction tuning (IT), including spoken based and text based tasks Alpaca 90 Multi -task ER + text based IT IEMOCAP Session 1-4, Alpaca 71 ASR + text based IT LibriSpeech,  Alpaca 274 ST + text based IT CoVoST2, MuST-C, Alpaca 343 SV + text based IT VoxCeleb, Alpaca 243 SQA + text based IT AMI, Fisher,Switchboard, Alpaca 773 ASR + ST LibriSpeech 74 ASR + SQA LibriSpeech 43 ASR + ST + text-based IT CoVoST2, Alpaca 5 ASR + SQA + text-based IT LibriSpeech, Alpaca 1066
Appendix G Implementation Details
As mentioned above, the semantic and acoustic speech encoders are the encoder of Whisper-large-v2888https://huggingface.co/openai/whisper-large-v2 and WavLM-base [^9] https://huggingface.co/microsoft/wavlm-base, the backbone LLM is LLaMA-2-chat-7b [^10] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and all of their parameters are frozen. The outputs of both modality adapters have a time stride of 80 ms and a dimension of 2048, and the rank ($R$) of LoRA is set as 32. In the first mixed single-task training stage, the total number of parameters in our model is 7.55 billion, of which 76.6 million are trainable. In the advanced training phase, the bottleneck dimension ($K$) of the prompt adapter is set as 1024. The 4096-dimensional prompt-dependent parameters produced by prompt adapter are element-wise multiplied with the outputs of the LoRA. Our models are trained with the two-stage curriculum-learning method on 32 V100 GPUs using the Adam optimizer, set with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.98$ and batch size equivalent to 30 seconds per GPU, where the first stage consisted of 400K steps and the subsequent stage involved an additional 150K steps. Additionally, we employed a maximum learning rate of $1\times 10^{-4}$, incorporating a warm-up phase for the first 10% of steps. The two-stage training data are presented in data construction part of Section 3.2.
Appendix H The Effect of Advanced Training for Single-tasks
Performance of model with or without advanced training on single-task instructions. mixed training means the first mixed single-task training stage, and advanced training means the second advanced multi-task training stage.
       Models ASR ST (En2De) SV ER SQQA SQA
test-clean test-other CoVoST2 MUSTC
WER BLEU Acc. Acc. Acc. Acc.
mixed training 2.0 4.8 23.9 21.9 0.91 0.72 0.55 67.30%
+ advanced training 2.0 4.8 23.6 21.7 0.91 0.72 0.57 67.55%

### v3

@@ 5 Conclusion
In this paper, we propose WavLLM, a robust and adaptive speech large language model, which uses LLaMA-2-chat as foundational LLM backbone, and extracts semantic and acoustic information from speech audio utilizing Whisper and WavLM encoders. Utilizing a curriculum learning approach, the proposed WavLLM commences with single-task instruction training in the initial mixed training phase and subsequently expanding our training to encompass additional multi-task prompts in the second advanced phase with the integration of the proposed prompt adapter. Massive experiments demonstrate that our WavLLM model delivers state-of-the-art performance on various speech-related tasks and robust generalization abilities on single-task and multi-task instructions, enhanced by a CoT processing ability that greatly improves its effectiveness in tackling intricate tasks.
Limitations
Although WavLLM shows a remarkable proficiency in handling speech-related tasks and impressive cross-modal instruction following and generalization capacity, it also exhibits some specific constraints.
Adaptive Use of CoT
Our WavLLM model produces performance improvements using CoT based instructions compared to one-shot non-CoT based instructions. However, it lacks the capability to autonomously decompose complex one-shot non-CoT based tasks into CoT based ones. For future work, we are interested in advancing the capability of adaptive use of CoT. This requires WavLLM to determine whether a task can be decomposed into multiple sub-tasks, and then applying the CoT approach accordingly.
Broader Applicability
Although our WavLLM model focuses primarily on speech in English, it can be readily extended to accommodate multiple languages. Additionally, the WavLLM model excels at processing and comprehending spoken language, yet it lacks the capability to generate speech. We defer the task of expanding WavLLM’s capabilities to synthesize speech to future research.
Safety and Ethics
Employing continuous speech representations within our WavLLM model may render it more vulnerable to adversarial attacks, potentially undermining the model’s compliance with the HHN criteria (Harmless, Helpful, Honest). This vulnerability merits further investigation for solutions.